One of two complementary angles is 2 less than three times the other angle. Find the larger angle.
Answer by nerdybill(7384)

One of two complementary angles is 2 less than three times the other angle. Find the larger angle.
.
Two angles are complementary when the sum is 90 degrees.
.
Let x= one of two angles
then
3x-2 = second angle
.
x + 3x-2 = 90
4x - 2 = 90
4x = 92
x = 23 degrees
.
Second angle:
3x-2 = 3(23)-2 = 69-2 = 67 degrees
.
Answer: 67 degrees